Suppose you are going to have pizza with several friends. You are thinking of getting one large pizza - say its diameter is 16 inches. However, some people are vegetarians, and so you decide to get two little pizzas, each of which is 8 inches in diameter. Suddenly, someone asks, "Are we getting the same amount of pizza if we get two little pizzas instead of one large one?"
What do you think? Justify your response.
Can someone explain this to me? Thanks

Answer by narayaba(40)

The answer is we are not getting the same amount of pizza
If you get a large one. The area of of the pizza is pi*(r^2) where r is radius
for 16inche pizza the radius is 16/2 =8 therefore area =pi*(8^2) = 64*pi
if you find the area of one small pizza with radius 8/2 =4 inches is pi*(4^2)
The area of two small pizzas is 2*pi*(4^2) = 1*16*pi =32*pi
64*pi > 32*pi
The amount we get by buying two little pizzas is half less than what we would get if we were to buy a large pizza.
32*pi = 0.5*64*pi